# A Survey of Frontiers in LLM Reasoning: Inference Scaling, Learning to Reason, and Agentic Systems

### v3

@@ Abstract
Figure 1: The LLM reasoning surge. We show the cumulative number (in thousands) of papers published from 2022 to 2/2025, based on Semantic Scholar keyword search. Research on reasoning regimes and agent architectures has accelerated notably since the introduction of Chain-of-Thought (CoT) in 2022. This growth is further influenced by other major developments, such as the release of ChatGPT (Ouyang et al., 2022) in 9/2022, and popularity of in-context learning (Brown et al., 2020) as an inference-time optimization method.
@@ 1 Introduction
Reasoning is the cognitive process of analyzing evidence, constructing arguments, and applying logic to form conclusions or make informed judgments. It is essential to many intellectual pursuits, including decision-making, problem-solving, and critical thinking. The study of reasoning spans multiple disciplines—philosophy (Passmore, 1961), psychology (Wason & JohnsonLaird, 1972), and computer science (Huth & Ryan, 2004)—as it provides insights into how individuals interpret information, evaluate alternatives, and develop sound conclusions using logic.
Recently, large language models (LLMs) have demonstrated a range of emerging abilities, such as in-context learning (Dong et al., 2024), role playing (Shanahan et al., 2023b) and domain adaptation (Ke et al., 2023; 2025a; Ke & Liu, 2023) as they scale, with reasoning becoming one of the most critical capabilities. As shown in Figure 1, this area has rapidly gained research attention, often referred to as LLM reasoning or reasoning language model (RLM) (Besta et al., 2025). The increasing focus on this topic is understandable, as reasoning capability is: (i) Challenging, requiring multi-step processing beyond the token-by-token generative nature of auto-regressive LLMs; (ii) Fundamental, as it is a core aspect of intelligence, particularly in planning and strategic decision-making; and, most importantly, (iii) Promising, as recent advances in LLMs hint at a viable path forward. Given these factors, reasoning is widely regarded as a prerequisite for more advanced AI systems approaching Artificial General Intelligence (AGI), beyond the conventional AI that aims to closely follow instruction (Duenas & Ruiz, 2024).
@@ 1.1 Reasoning Regimes
Several variants of prompting methods (Paranjape et al., 2021; Sanh et al., 2022; Mishra et al., 2022) have been introduced, providing structured prompts to enhance reasoning. Additionally, inference scaling optimizes reasoning through search and planning (Dua et al., 2022; Zhou et al., 2023a; Khot et al., 2023; Suzgun & Kalai, 2024a). One key challenge in search and planning is evaluating the quality of candidate solutions. However, evaluating reasoning quality is inherently difficult, even for humans. Existing approaches can be categorized based on whether they judge the final outcome, i.e., outcome reward models (ORMs) (Hendrycks et al., 2021b), or the reasoning process, i.e., process reward models (PRMs) (Lightman et al., 2024).
@@ 1.4 Goal and Structure of the Survey
The goal of this survey is to provide a comprehensive overview of key algorithmic details and major milestones in LLM reasoning research, particularly since the emergence of Chain-of-Thought (CoT), across both regime and architecture dimensions. We believe this is a timely and valuable contribution to the community, given the clear acceleration in research following CoT’s introduction in 2022 (Figure 1). The rapid growth in studies exploring all aspects of LLM reasoning—from regimes and architectures to training algorithms—highlights the increasing importance and utility of reasoning capabilities in advancing the field.
@@
[…]
Open
Challenges §6, draw=bluishgreen, fill=bluishgreen!15, text width=1.8cm, align=left [Understand
Reasoning §6.2.2, draw=bluishgreen, fill=bluishgreen!10, text width=2.2cm, align=left [Evaluation: Hendrycks et al. (2021b); Hao et al. (2024a); Wang & Zhao (2024), draw=bluishgreen, fill=bluishgreen!5, text width=11.2cm, align=left] [Formal Analysis: Han et al. (2022); Li et al. (2025b); Saparov & He (2023), draw=bluishgreen, fill=bluishgreen!5, text width=11.2cm, align=left] [Empirical Analysis: Wu et al. (2024a); Dutta et al. (2024); Sprague et al. (2024a), draw=bluishgreen, fill=bluishgreen!5, text width=11.2cm, align=left] ] [Domain-specific
Reasoning §6.1.3, draw=bluishgreen, fill=bluishgreen!10, text width=2.2cm, align=left [Game Reasoning: Lee et al. (2024); Guo et al. (2023); Yang et al. (2024e), draw=bluishgreen, fill=bluishgreen!5, text width=11.2cm, align=left] [Tabular Reasoning: Chen (2023); Ye et al. (2023b); Sui et al. (2024), draw=bluishgreen, fill=bluishgreen!5, text width=11.2cm, align=left] [Code Generation: OpenAI et al. (2025); Team et al. (2025); Yang et al. (2024a), draw=bluishgreen, fill=bluishgreen!5, text width=11.2cm, align=left] [Mathematical Reasoning: Yu et al. (2024b); Shao et al. (2024); Jiang et al. (2024a), draw=bluishgreen, fill=bluishgreen!5, text width=11.2cm, align=left] ] ] [Learning to
Reason §5, draw=vermillion, fill=vermillion!15, text width=1.8cm, align=left [Inf & cost-aware
Learning §5.4, draw=vermillion, fill=vermillion!10, text width=2.2cm, align=left [Inference-aware Training: Balashankar et al. (2024); Chow et al. (2024), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] [Cost-aware Training: Damani et al. (2025); Snell et al. (2024); Ye et al. (2025), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] ] [Multi-agent
System §5.3, draw=vermillion, fill=vermillion!10, text width=2.2cm, align=left [Action Coordination: Lau et al. (2012); Xu et al. (2023); Li et al. (2024f), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] [Agent-Agent Communication: Zhang et al. (2021); Hong et al. (2023), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] ] [Single-agent
System §5.2, draw=vermillion, fill=vermillion!10, text width=2.2cm, align=left [Reinforcement Learning: Shao et al. (2024); Havrilla et al. (2024), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] [Preference Learning: Rafailov et al. (2023); Jiao et al. (2024a); Lai et al. (2024), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] [Agent-Environment Interaction: Gou et al. (2024); Yin et al. (2024), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] ] [Standalone
LLM §5.1, draw=vermillion, fill=vermillion!10, text width=2.2cm, align=left [Training from Trajectories: Huang et al. (2024d); Wang et al. (2023d), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] [Trajectory Collection: Zelikman et al. (2022); Yuan et al. (2023); Qin et al. (2024), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] [Prompt Construction: Xu et al. (2024a); Yu et al. (2024c); Li et al. (2024a), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] ] ] [Learning
Algorithms §4, draw=reddishpurple, fill=reddishpurple!15, text width=1.8cm, align=left [Learning of
Verifiers §4.2, draw=reddishpurple, fill=reddishpurple!10, text width=2.2cm, align=left [Generative Verifiers: Yao et al. (2023c); Zhang et al. (2024f); Mahan et al. (2024), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] [Process Reward Models: Wang et al. (2024g); Jiao et al. (2024a); Setlur et al. (2024b), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] [Outcome Reward Models: Liu et al. (2024a), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] ] [Learning of
Reasoner §4.1, draw=reddishpurple, fill=reddishpurple!10, text width=2.2cm, align=left [Preference Learning: Rafailov et al. (2023); Ethayarajh et al. (2024), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] [Reinforcement Learning: Schulman et al. (2017); Shao et al. (2024), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] ] ] [Inference
Scaling §3, draw=skyblue, fill=skyblue!15, text width=1.8cm, align=left [Multi-agent
System §3.3, draw=skyblue, fill=skyblue!10, text width=2.2cm, align=left [Action Coordination: Wang et al. (2024c); Gao et al. (2024b); Liu et al. (2024b), draw=skyblue, fill=skyblue!5, text width=11.2cm, align=left] [Communication Design: Liang et al. (2023b); Chen et al. (2023c); Liu et al. (2024e), draw=skyblue, fill=skyblue!5, text width=11.2cm, align=left] ] [Single-agent
System §3.2, draw=skyblue, fill=skyblue!10, text width=2.2cm, align=left [Retrieval and Tool Use: Lu et al. (2024a); Hammane et al. (2024); Ma et al. (2024b), draw=skyblue, fill=skyblue!5, text width=11.2cm, align=left] [Verifier and Reflection: First et al. (2023); Xin et al. (2024a); Tian et al. (2024), draw=skyblue, fill=skyblue!5, text width=11.2cm, align=left] ] [Standalone
LLM §3.1, draw=skyblue, fill=skyblue!10, text width=2.2cm, align=left [Search & Planning: Dua et al. (2022); Yao et al. (2023a); Besta et al. (2024), draw=skyblue, fill=skyblue!5, text width=11.2cm, align=left] [Prompt Engineering: Zhou et al. (2023b); Wei et al. (2022b); Yasunaga et al. (2024), draw=skyblue, fill=skyblue!5, text width=11.2cm, align=left] ] ] ]
@@ 1.5 Comparison to Related Surveys
Reasoning in LLMs has long been a fundamental challenge in the field. Earlier works, such as Huang & Chang (2023), provide a comprehensive overview of the evolution of informal deductive reasoning covering developments prior to the emergence of LLM agents and Reasoning Language Models (RLMs). Our work extends this discussion by focusing on LLM agents and RLMs. Qiao et al. (2023b) offer a detailed summary of advancements in LLM reasoning, with a particular emphasis on prompting techniques. In contrast, we offer a broader range of regimes (from inference to training) and architectures (from standalone LLM to multi-agent systems). Readers interested in a formal definition and taxonomy of natural language reasoning—grounded in philosophical foundations—may refer to Yu et al. (2024a), which focuses specifically on this direction and is complementary to our scope.
@@ 2.1 Problem Formulation
LLM reasoning is often formulated within the Markov Decision Process (MDP) framework (Bellman, 1958), treating reasoning as a sequential decision-making process. While many of the terminologies in LLM reasoning originate from the AI agent and reinforcement learning (RL) literature (Russell & Norvig, 2010), their meaning in LLM reasoning can sometimes differ to suit the nature of LLM-based reasoning.
@@ 2.3.2 From Standalone LLM to Language Agents
While the concept of an agent has been a long-standing idea in AI (Russell & Norvig, 2010), the notion of language agents has gained prominence alongside recent advancements in LLMs. [^3] In this survey, the terms agent and LLM-based agent are used interchangeably unless stated otherwise. The key distinction between an agent and a standalone LLM lies in two advanced capabilities: interactiveness (Weng, 2023; Yao & Narasimhan, 2023) and autonomy (Xi et al., 2023; Wang et al., 2024d). Interactiveness refers to an agent’s ability to engage with the external world, including environments or other agents. This capability is crucial because LLMs, while powerful, often have limited knowledge and reasoning abilities confined to their internal memory. By enabling interaction with the external world, an LLM can augment its internal knowledge with external information, significantly expanding its understanding and grounding its outputs in real-world observations. Autonomy, on the other hand, refers to an agent’s ability not only to follow human instructions but also to independently initiate and execute actions. This capability often involves planning but can extend to more complex behaviors. For instance, a fully autonomous agent should be capable of detecting novel situations, proactively taking initiative, and determining effective interaction strategies without explicit human guidance. These advanced capabilities distinguish LLM-based agents from standalone LLMs, enabling them to operate more dynamically and adaptively in real-world scenarios.
@@ 2.3.3 Single-agent Systems
While a fully autonomous agent should ideally learn to interact with the environment automatically, the literature identifies several predefined interaction patterns (also referred to as workflows (Schluntz & Zhang, 2024)) that have proven effective. We elaborate on these patterns below and, in Sections 3.2 and 5.2, explore specific techniques that leverage them to improve agent performance.
@@ 3 Improving Reasoning with Inference Scaling
Constructing
Prompts
Instruction engineering Modify instruction by human-design template Paranjape et al. (2021); Zhou et al. (2023b)
@@
Optimizing
Output
Generating subtasks Decompose the original task into manageable subtasks Dua et al. (2022); Zhou et al. (2023a)
@@ 3.1 Inference Scaling With Standalone LLM
In this section, we examine the core components and techniques that have made inference-time reasoning methods effective. Many of these methods draw inspiration from research on human cognitive processes on planning, problem solving, and decision-making (Newell et al., 1959; 1972; Stanovich & West, 2000).
@@ 3.1.2 Optimizing Reasoning Output with Search and Planning
Simple CoT prompting often struggles as task complexity increases, particularly when the task surpasses the complexity of the provided demonstrations. To address this, researchers have proposed methods that explicitly guide models in decomposing tasks into subtasks, thereby enhancing intermediate step reasoning. Dua et al. (2022) propose an iterative approach, where tasks are progressively broken down into simpler subtasks and solved step-by-step. Similarly, Zhou et al. (2023a); Khot et al. (2023) and Suzgun & Kalai (2024a) advocate for a “divide-and-conquer” strategy, where tasks are first divided into subtasks and then solved sequentially.
@@
Research on human problem-solving reveals that complex reasoning tasks often admit multiple valid paths to reach a correct solution (Stanovich & West, 2000). Compared to linear reasoning structures like chain-of-thought, approaches that incorporate exploration during problem-solving have shown significant improvements for complex reasoning tasks. Unlike task decomposition methods (Dua et al., 2022; Zhou et al., 2023a; Khot et al., 2023), exploration-based approaches employ dynamic search through multiple possible reasoning paths simultaneously rather than following certain decomposition patterns, enabling models to explore ambiguous solution strategies for complex problems. Exploration typically involves two key components: branching and aggregation. Due to the stochastic nature of language model decoding, branching is often implemented through independent re-sampling with non-zero temperature, generating diverse reasoning chains. Early methods, such as self-consistency (Wang et al., 2023f), introduced branching only at the beginning of the reasoning chain, conditioned on the initial query. While simple, this approach lacks local exploration of intermediate reasoning steps, has limited applicability for tasks with multiple valid answers, and produces reasoning chains with restricted diversity (Chen et al., 2024d). More recent advancements, such as Tree-of-Thoughts (Yao et al., 2023a), Graph-of-Thoughts (Besta et al., 2024), and Forest-of-Thoughts (Bi et al., 2024), enable finer-grained branching by considering both the query and a history of previous thoughts or thought-state sequences, allowing for more nuanced and flexible exploration.
The effectiveness of branched reasoning paths with thoughts or answers depends on aggregation or evaluation strategies. Recent progress is centered around two categories: ensemble-based methods and verifier-based methods. Ensemble-based methods have been widely employed due to their simplicity and self-contained nature, requiring no external knowledge or sources for validation. These approaches typically employ strategies such as majority voting across answer tokens (Wang et al., 2023f; 2024a; Li et al., 2024b) or confidence-based selection (Wang & Zhou, 2024). Verifier-based methods, in contrast, employ external verifiers or judges to score and select preferred answers among candidate solutions.
@@
Action Enhancement Retrieval and Tool Access external knowledge and specialized resources Li et al. (2024e); Ma et al. (2024a)

### v4

@@ Abstract
Figure 1: The LLM reasoning surge. We show the cumulative number (in thousands) of papers published from 2022 to 2/2025, based on Semantic Scholar keyword search. Research on reasoning regimes and agent architectures has accelerated notably since the introduction of Chain-of-Thought (CoT) in 2022. This growth is further influenced by other major developments, such as the release of ChatGPT (Ouyang et al., 2022) in 9/2022, and popularity of in-context learning (brown2020languagemodelsfewshotlearners) as an inference-time optimization method.
@@ 1 Introduction
Reasoning is the cognitive process of analyzing evidence, constructing arguments, and applying logic to form conclusions or make informed judgments. It is essential to many intellectual pursuits, including decision-making, problem-solving, and critical thinking. The study of reasoning spans multiple disciplines—philosophy (Passmore, 1961), psychology (Wason and JohnsonLaird, 1972), and computer science (Huth and Ryan, 2004)—as it provides insights into how individuals interpret information, evaluate alternatives, and develop sound conclusions using logic.
Recently, large language models (LLMs) have demonstrated a range of emerging abilities, such as in-context learning (Dong et al., 2024), role playing (Shanahan et al., 2023b) and domain adaptation (Ke et al., 2023; 2025a; Ke and Liu, 2023) as they scale, with reasoning becoming one of the most critical capabilities. As shown in Figure LABEL:fig.trend, this area has rapidly gained research attention, often referred to as LLM reasoning or reasoning language model (RLM) (Besta et al., 2025). The increasing focus on this topic is understandable, as reasoning capability is: (i) Challenging, requiring multi-step processing beyond the token-by-token generative nature of auto-regressive LLMs; (ii) Fundamental, as it is a core aspect of intelligence, particularly in planning and strategic decision-making; and, most importantly, (iii) Promising, as recent advances in LLMs hint at a viable path forward. Given these factors, reasoning is widely regarded as a prerequisite for more advanced AI systems approaching Artificial General Intelligence (AGI), beyond the conventional AI that aims to closely follow instruction (Duenas and Ruiz, 2024).
@@ 1.1 Reasoning Regimes
Several variants of prompting methods (Paranjape et al., 2021; Sanh et al., 2022; Mishra et al., 2022) have been introduced, providing structured prompts to enhance reasoning. Additionally, inference scaling optimizes reasoning through search and planning (Dua et al., 2022; Zhou et al., 2023a; Khot et al., 2023; Suzgun and Kalai, 2024a). One key challenge in search and planning is evaluating the quality of candidate solutions. However, evaluating reasoning quality is inherently difficult, even for humans. Existing approaches can be categorized based on whether they judge the final outcome, i.e., outcome reward models (ORMs) (Hendrycks et al., 2021b), or the reasoning process, i.e., process reward models (PRMs) (Lightman et al., 2024).
@@ 1.4 Goal and Structure of the Survey
The goal of this survey is to provide a comprehensive overview of key algorithmic details and major milestones in LLM reasoning research, particularly since the emergence of Chain-of-Thought (CoT), across both regime and architecture dimensions. We believe this is a timely and valuable contribution to the community, given the clear acceleration in research following CoT’s introduction in 2022 (Figure LABEL:fig.trend). The rapid growth in studies exploring all aspects of LLM reasoning—from regimes and architectures to training algorithms—highlights the increasing importance and utility of reasoning capabilities in advancing the field.
@@ 1.5 Comparison to Related Surveys
Reasoning in LLMs has long been a fundamental challenge in the field. Earlier works, such as Huang and Chang (2023), provide a comprehensive overview of the evolution of informal deductive reasoning covering developments prior to the emergence of LLM agents and Reasoning Language Models (RLMs). Our work extends this discussion by focusing on LLM agents and RLMs. Qiao et al. (2023b) offer a detailed summary of advancements in LLM reasoning, with a particular emphasis on prompting techniques. In contrast, we offer a broader range of regimes (from inference to training) and architectures (from standalone LLM to multi-agent systems). Readers interested in a formal definition and taxonomy of natural language reasoning—grounded in philosophical foundations—may refer to Yu et al. (2024a), which focuses specifically on this direction and is complementary to our scope.
@@ 2.1 Problem Formulation
LLM reasoning is often formulated within the Markov Decision Process (MDP) framework (Bellman, 1958), treating reasoning as a sequential decision-making process. While many of the terminologies in LLM reasoning originate from the AI agent and reinforcement learning (RL) literature (Russell and Norvig, 2010), their meaning in LLM reasoning can sometimes differ to suit the nature of LLM-based reasoning.
@@ 2.3.2 From Standalone LLM to Language Agents
While the concept of an agent has been a long-standing idea in AI (Russell and Norvig, 2010), the notion of language agents has gained prominence alongside recent advancements in LLMs. [^3] In this survey, the terms agent and LLM-based agent are used interchangeably unless stated otherwise. The key distinction between an agent and a standalone LLM lies in two advanced capabilities: interactiveness (Weng, 2023; Yao and Narasimhan, 2023) and autonomy (Xi et al., 2023; Wang et al., 2024d). Interactiveness refers to an agent’s ability to engage with the external world, including environments or other agents. This capability is crucial because LLMs, while powerful, often have limited knowledge and reasoning abilities confined to their internal memory. By enabling interaction with the external world, an LLM can augment its internal knowledge with external information, significantly expanding its understanding and grounding its outputs in real-world observations. Autonomy, on the other hand, refers to an agent’s ability not only to follow human instructions but also to independently initiate and execute actions. This capability often involves planning but can extend to more complex behaviors. For instance, a fully autonomous agent should be capable of detecting novel situations, proactively taking initiative, and determining effective interaction strategies without explicit human guidance. These advanced capabilities distinguish LLM-based agents from standalone LLMs, enabling them to operate more dynamically and adaptively in real-world scenarios.
@@ 2.3.3 Single-agent Systems
While a fully autonomous agent should ideally learn to interact with the environment automatically, the literature identifies several predefined interaction patterns (also referred to as workflows (Schluntz and Zhang, 2024)) that have proven effective. We elaborate on these patterns below and, in Sections 3.2 and 5.2, explore specific techniques that leverage them to improve agent performance.
@@ 3 Improving Reasoning with Inference Scaling
Constructing Prompts Instruction engineering Modify instruction by human-design template Paranjape et al. (2021); Zhou et al. (2023b)
@@
Optimizing Output Generating subtasks Decompose the original task into manageable subtasks Dua et al. (2022); Zhou et al. (2023a)
@@ 3.1 Inference Scaling With Standalone LLM
In this section, we examine the core components and techniques that have made inference-time reasoning methods effective. Many of these methods draw inspiration from research on human cognitive processes on planning, problem solving, and decision-making (Newell et al., 1959; 1972; Stanovich and West, 2000).
@@ 3.1.2 Optimizing Reasoning Output with Search and Planning
Simple CoT prompting often struggles as task complexity increases, particularly when the task surpasses the complexity of the provided demonstrations. To address this, researchers have proposed methods that explicitly guide models in decomposing tasks into subtasks, thereby enhancing intermediate step reasoning. Dua et al. (2022) propose an iterative approach, where tasks are progressively broken down into simpler subtasks and solved step-by-step. Similarly, Zhou et al. (2023a); Khot et al. (2023) and Suzgun and Kalai (2024a) advocate for a “divide-and-conquer” strategy, where tasks are first divided into subtasks and then solved sequentially.
@@
Research on human problem-solving reveals that complex reasoning tasks often admit multiple valid paths to reach a correct solution (Stanovich and West, 2000). Compared to linear reasoning structures like chain-of-thought, approaches that incorporate exploration during problem-solving have shown significant improvements for complex reasoning tasks. Unlike task decomposition methods (Dua et al., 2022; Zhou et al., 2023a; Khot et al., 2023), exploration-based approaches employ dynamic search through multiple possible reasoning paths simultaneously rather than following certain decomposition patterns, enabling models to explore ambiguous solution strategies for complex problems. Exploration typically involves two key components: branching and aggregation. Due to the stochastic nature of language model decoding, branching is often implemented through independent re-sampling with non-zero temperature, generating diverse reasoning chains. Early methods, such as self-consistency (Wang et al., 2023f), introduced branching only at the beginning of the reasoning chain, conditioned on the initial query. While simple, this approach lacks local exploration of intermediate reasoning steps, has limited applicability for tasks with multiple valid answers, and produces reasoning chains with restricted diversity (Chen et al., 2024d). More recent advancements, such as Tree-of-Thoughts (Yao et al., 2023a), Graph-of-Thoughts (Besta et al., 2024), and Forest-of-Thoughts (Bi et al., 2024), enable finer-grained branching by considering both the query and a history of previous thoughts or thought-state sequences, allowing for more nuanced and flexible exploration.
The effectiveness of branched reasoning paths with thoughts or answers depends on aggregation or evaluation strategies. Recent progress is centered around two categories: ensemble-based methods and verifier-based methods. Ensemble-based methods have been widely employed due to their simplicity and self-contained nature, requiring no external knowledge or sources for validation. These approaches typically employ strategies such as majority voting across answer tokens (Wang et al., 2023f; 2024a; Li et al., 2024b) or confidence-based selection (Wang and Zhou, 2024). Verifier-based methods, in contrast, employ external verifiers or judges to score and select preferred answers among candidate solutions.
@@
Action Enhancement Retrieval and Tool Access external knowledge and specialized resources Li et al. (2024e); Ma et al. (2024b)